solve dx/(x-y)y^2 = dy/(y-x)x^2 = dz/x^2+y^2


1
Expert's answer
2021-04-29T16:47:27-0400




The question is

"\bigstar"

"dx\over { ((x-y)y^2) }" ="dy\over((y-x)x^2)" = "dz\over(x^2+y^2)"



Consider these



"dx\over { ((x-y)y^2) }" ="dy\over((y-x)x^2)"


We get

Integration of

"\bull"

"\boxed{\int y^2dy = \int -x^2dx}"



"\bull"

"\boxed{y^3 = -x^3+c1}"




"\bigstar"

"{dx\over ((x-y)y^2)}={dz\over(x^2+y^2)}"



 "\implies"


"\boxed{ \int {y^2} dz= \int {(x^2+y^2)\over (x-y) }dx }"


"\implies"


"\boxed{∫ {(x^2+y^2)\over (x-y) }dx}" ="\boxed{∫ {2y^2\over(x-y)+x+y)}dx}" =




"\boxed{2y^2\intop{1\over(x-y)}dx +\intop x dx + y \intop 1 dx}="


= "\boxed{2 y^2 ln(x-y) + {x^2\over{2 }}+ xy + c_2}"



"\bull"

"\bigstar"

"\boxed{y^2ln(z) = 2 y^2 ln(x-y) + {x^2\over{2 }}+ xy + c_2

}"


integral surface:

"\bull"

"\boxed{ y^3 = -x^3+c_1 }"



"\bull"

"\boxed{y^2ln(z) = 2 y^2 ln(x-y) + {x^2\over{2 }}+ xy + c_2

}"